OpenStudy (anonymous):

MEDAL to the first correct answer ;) step by step explanation... (Using the quadratic formula) In a right angled triangle, the second to longest side is 5 cm longer than the shortest side, and the hypotenuse is three times longer than the shortest side. Find the exact length of the hypotenuse.

OpenStudy (anonymous):

Shortest side = x other side = x+5 hypotenuse = 3x we know, in a right angled triangle , (short side)^2 + (long side)^2 = (hypotenuse)^2 (x)^2 + (x+5)^2 = (3x)^2

OpenStudy (anonymous):

Yeah I know that :)

OpenStudy (anonymous):

But can you solve for x, please?

OpenStudy (anonymous):

Haha, So all you need is a confirmation, no ? :P

OpenStudy (anonymous):

I tried solving it but I got a wrong answer so I wanna know what I did wrong :p

OpenStudy (anonymous):

What answer di you get , anyways?

OpenStudy (anonymous):

answer*

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{(-b)^2 -4ac} }{ 2a }\] the lenghth is always positif so u'll leave out any negative answer

OpenStudy (anonymous):

it's not (-b) squared it's b squared

OpenStudy (anonymous):

check again ur simplification u had \[\frac{ 10 + 20\sqrt{2} }{ 14 }\]

OpenStudy (anonymous):

(-b)^2=(b)^2 isn't it?

OpenStudy (anonymous):

yeah but the formula is b squared and not -b squared even if it's the same answer...

OpenStudy (anonymous):

Look, I get \[\frac{5 \pm 10\sqrt2}{7}\]

OpenStudy (anonymous):

The thing that you have to see is that , the value that you get, i.e (5 + 10√2)/7 , is the value of x .. But as you can see, we assumes hypotenuse to be 3x , thus, you have to multiply 3 with this value of x .. such that you get : 3( (5 + 10√2)/7 ) = 15 + 30√2 / 7 .. Understood ? @IBstudent

OpenStudy (anonymous):

Oh OMG Thank you!!! now I got it!!

OpenStudy (anonymous):

You're welcome :D